# Laser: Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

### v2

@@ 5 Discussion and Conclusion
While our pipeline outperforms baselines in almost all settings, it shows only little improvements when sampling large portions from weak source data (as shown in Section 4.2.2). This might be due to a potential ceiling effect on possible performance given the weak source material, which might require improving the source data rather than elaborate sampling techniques. The modular setup of Laser allows for further incremental improvements of different scorers in future work.
@@
A major issue in collecting data for our difficulty scorer is a certain unreliability in the evaluation of model responses. Previous research shows that evaluations oftentimes have weak robustness to e.g. the prompt formatting (Weber et al., 2023a, b; Sclar et al., 2024; Polo et al., 2024b), bias of LLMs-as-a-judge (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024) or errors in post-processing (such as issues in extracting the answer from the model response). For example, GPT4o showed overall weaker performance on some of our evaluated subsets than some small open-source models. Upon closer inspection, we encountered that – while providing the correct answer – GPT4o generally tends not to follow the formatting of the given few-shot examples and rather responds in an open-form manner, resulting in failing the tight response search masks of the used evaluation frameworks. While we try to mitigate this issue as much as possible, we cannot guarantee that difficulty scores exactly reflect a model’s capacity to solve a given data point.
@@
Our results indicate that we get only small performance gains in setups with very weak source data. However, there might be a ceiling for possible performance in such settings, as data selection alone might not be sufficient to achieve good performance in such setups.
@@ Acknowledgments
We want to thank Rishiraj Saha Roy, Lucas Druart and Christian Kroos for their valuable feedback on earlier versions of this manuscript. This work was partly funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) and partly by the Free State of Bavaria in the DSgenAI project (Grant Nr.: RMF-SG20-3410-2-18-4). Further, the authors gratefully acknowledge the Centre for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] and ScaDS.AI at the Technical University of Dresden for providing its facilities for experimental work.
@@ A.2.1 Generating $\mathcal{D}_{\text{diff}}$
During the data collection for the difficulty scorer, we collect training sets from different benchmarks as well as data points from OpenAssistent (Köpf et al., 2023) as samples of open-ended generation. We evaluate these samples using the existing evaluation frameworks LM-evaluation harness (Gao et al., 2024), big-code evaluation harness (Ben Allal et al., 2022) and FastChat (Zheng et al., 2023b). For the LLM-as-a-judge evaluation, we use GPT4o as the judge and evaluate only a subset of 4 LLMs from Table 4 that we deem representative in terms of capabilities (gpt-4o-2024-08-06, Mistral-Small-24B-Instruct-2501, Mistral-7B-Instruct-v0.3 and SmolLM2-1.7B-Instruct).
@@ A.2.3 Effectiveness
To evaluate the effectiveness of the difficulty scorer for data selection across benchmarks, we sample 25k data points from the datasets in Table 2, either at random or based on Laser difficulty score ranking. The results in Figure 10(a) show that difficulty-based sampling outperforms random sampling on all benchmarks except GPQA. For comparison, Figure 10(b) reports results using the Deita complexity scorer under the same setup.
@@ A.5.2 Datasets robustness experiments
Dataset #samples (#turns) ConiferLM/Conifer (Sun et al., 2024) 13k databricks/databricks-dolly-15k (Conover et al., 2023) 15k akoksal/LongForm (Köksal et al., 2023) 23k alpaca (Taori et al., 2023) 52K vicgalle/alpaca-gpt4 52K ai2-adapt-dev/flan_v2_converted (Sanh et al., 2022) 90K nvidia/Daring-Anteater (Wang et al., 2024) 100k AI-MO/NuminaMath-CoT999random subset (Li et al., 2024a) 250k all 595K
@@ A.7 Results – details
OpenBookQA (OpenBookQA2018) Default train 299 acc
@@
MLQA (Lewis et al., 2019) 49 lang. combinations val 715 F1

### v3

@@ 5 Discussion and Conclusion
While our pipeline outperforms baselines in almost all settings, it shows only little improvements when sampling large portions from weak source data (as shown in Section 4.2.2). This may reflect a potential ceiling effect on possible performance given the weak source material, suggesting that improving the data itself—rather than relying solely on elaborate sampling techniques—might be necessary. In such cases, iterative data refinement seems promising, especially when guided by reliable scorers for assessing sample difficulty and quality. The modular setup of Laser allows for further incremental improvements of different scorers in future work.
@@
A major issue in collecting data for our difficulty scorer is a certain unreliability in the evaluation of model responses. Previous research shows that evaluations oftentimes have weak robustness to e.g. the prompt formatting (Weber et al., 2023a, b; Sclar et al., 2024; Polo et al., 2024b), bias of LLMs-as-a-judge (Panickssery et al., 2024; Stureborg et al., 2024; Wataoka et al., 2024) or errors in post-processing (such as issues in extracting the answer from the model response). For example, GPT4o showed overall weaker performance on some of our evaluated subsets than some small open-source models. Upon closer inspection, we encountered that—while providing the correct answer—GPT4o generally tends not to follow the formatting of the given few-shot examples and rather responds in an open-form manner, resulting in failing the tight response search masks of the used evaluation frameworks. While we try to mitigate this issue as much as possible, we cannot guarantee that difficulty scores exactly reflect a model’s capacity to solve a given data point.
@@
Our results indicate modest performance gains in setups with very weak source data. However, there might be a ceiling for possible performance in such settings, where data selection alone might not be sufficient to achieve good performance. In such cases, iterative data refinement appears to be a promising approach, particularly when supported by reliable scorers for assessing sample difficulty and quality.
@@ Acknowledgments
We thank Rishiraj Saha Roy, Lucas Druart and Christian Kroos for their valuable feedback on earlier versions of this manuscript. This work was supported in part by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the OpenGPT-X project (no. 68GX21007D) and in part by the Free State of Bavaria in the DSgenAI project (Grant Nr.: RMF-SG20-3410-2-18-4). We also gratefully acknowledge the Centre for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] and ScaDS.AI at the Technical University of Dresden for providing its facilities for experimental work.
@@ A.2.1 Generating $\mathcal{D}_{\text{diff}}$
During the data collection for the difficulty scorer, we collect training sets from different benchmarks as well as data points from OpenAssistent (Köpf et al., 2023) as samples of open-ended generation. We evaluate these samples using the existing evaluation frameworks LM-evaluation harness (Gao et al., 2024), big-code evaluation harness (Ben Allal et al., 2022) and FastChat (Zheng et al., 2023). For the LLM-as-a-judge evaluation, we use GPT4o as the judge and evaluate only a subset of 4 LLMs from Table 4 that we deem representative in terms of capabilities (gpt-4o-2024-08-06, Mistral-Small-24B-Instruct-2501, Mistral-7B-Instruct-v0.3 and SmolLM2-1.7B-Instruct).
@@ A.2.3 Effectiveness
To evaluate the effectiveness of the difficulty scorer for data selection across benchmarks, we sample 25k data points from the datasets in Table 2, either at random or based on Laser difficulty score ranking. The results in Figure 10(a) show that difficulty-based sampling outperforms or is comparable to random sampling on all benchmarks except GPQA. For comparison, Figure 10(b) reports results using the Deita complexity scorer under the same setup.
@@ A.5.2 Datasets robustness experiments
Dataset #samples (#turns) ConiferLM/Conifer (Sun et al., 2024) 13k databricks/databricks-dolly-15k (Conover et al., 2023) 15k akoksal/LongForm (Köksal et al., 2023) 23k alpaca (Taori et al., 2023) 52K vicgalle/alpaca-gpt4 52K ai2-adapt-dev/flan_v2_converted (Sanh et al., 2022) 90K nvidia/Daring-Anteater (Wang et al., 2024) 100k AI-MO/NuminaMath-CoT101010random subset (Li et al., 2024a) 250k all 595K
@@ A.7 Results – details
OpenBookQA (Mihaylov et al., 2018) Default train 299 acc
@@
MLQA (Lewis et al., 2020) 49 lang. combinations val 715 F1